future value of ordinary Annuity


metrobank is paying 8% interest compounded quarterly. Find the future value 10,000 deposited at the end of every three months for 6 years


1
Expert's answer
2021-03-22T15:50:40-0400
"FV=\dfrac{(1+i)^n-1}{i}\cdot R"

Given

"R=10000"

"r=\dfrac{0.08}{4}=0.02"

"n=4\cdot6=24"


"FV=\dfrac{(1+0.02)^{24}-1}{0.02}\cdot 10000=304218.62"

"FV=304218.62"